# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 10 Comparison with étale cohomology
We have $\mathrm{H}^{i}\bigl(X_{\mathrm{pket}},\widehat{\mathcal{L}}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}_{\mathrm{et}}\bigl(U,\mathcal{L}(r)\bigr)$. Similarly we have $\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U,\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\mathrm{pket}},j_{!}\bigl(\widehat{\mathcal{L}}(r)\bigr)\bigr)$ where, following [LLZ, Def. 2.3.2], we define $\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U,\mathcal{L}(r)\bigr):=\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},j_{!}\bigl(\mathcal{L}(r)\bigr)\bigr)$.
@@
We start with the first claim. First we have isomorphisms if we replace $\mathcal{L}$ with $\mathcal{L}/p^{n}\mathcal{L}$ for the isomorphisms $\mathrm{H}^{i}\bigl(X_{\mathrm{pket}},\widehat{\mathcal{L}}(r)/p^{n}\widehat{\mathcal{L}}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ by [DLLZa, Prop. 5.1.7] and $\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}_{\mathrm{et}}\bigl(U,\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ by [DLLZa, Cor. 4.6.7]. They are also finite modules ; thanks to [DLLZa, Thm. 6.2.1] the groups $\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ are finite and vanish for $i>2\mathrm{dim}X$. Using the Hochschild-Serre spectral sequence this implies that also $\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ is finite for every $i$. In particular, the inverse system for varying $n$ is Miitag-Leffler and the result follows.
For the second claim one argues similarly, using the finiteness and the vanishing for $i>2\mathrm{dim}X$ of the group $\mathrm{H}^{i}_{c}\bigl(X_{\overline{K},\mathrm{ket}},j_{!}\bigl(\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)\bigr)$ which is the content of [LLZ, Thm. 2.3.5]. The isomorphism $\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},j_{\ast}(\mathcal{L}(r))\bigr)$ and last statements for the de Rham period sheaves follow from [DLLZb, Lemma 3.6.1] using [DLLZa, Thm 6.2.1 & Cor. 6.3.4]. For the crystalline statement it suffices to show that the natural map $\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},j_{\ast}(\mathcal{L}(r))\bigr)\otimes A_{\mathrm{crys}}\to\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{A}_{\mathrm{crys}}\bigr)$ is an isomorphism. Since $p$ is a regular element in $\mathbb{A}_{\mathrm{crys}}$ by [AI, Cor. 3.27], we are reduced to prove that the natural map $\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},j_{\ast}(\mathcal{L}(r))\bigr)\otimes A_{\mathrm{crys}}/pA_{\mathrm{crys}}\to\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{A}_{\mathrm{crys}}/p\mathbb{A}_{\mathrm{crys}}\bigr)$ is an isomorphism. As $\mathbb{A}_{\mathrm{crys}}/p\mathbb{A}_{\mathrm{crys}}$ is a free module over $\mathcal{O}_{X_{\mathrm{pket}}}^{+}/p\mathcal{O}_{X_{\mathrm{pket}}}^{+}$ and $A_{\mathrm{crys}}/pA_{\mathrm{crys}}$ is a free $\mathcal{O}_{\overline{K}}/p\mathcal{O}_{\overline{K}}$ with the same basis by [AI, Cor. 3.24], one is reduced to the primitive comparison theorem [DLLZa, Thm 6.2.1].
@@
$0\longrightarrow j_{!}\bigl(\widehat{\mathcal{L}}(r)|_{U}\bigr)[p^{-1}]\longrightarrow\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\stackrel{{\scriptstyle\iota\circ(1-\frac{\varphi}{p^{r}}),\iota)}}{{\longrightarrow}}\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\oplus\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathrm{Gr}^{r}\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}\longrightarrow 0.$
@@
The first two sequences are obtained from the analogous sequence with $\mathcal{L}=\mathbb{Z}_{p}$. Hence, we can assume that $\mathcal{L}=\mathbb{Z}_{p}$. Since $\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{crys}}=t^{r}\mathrm{Fil}^{0}\mathbb{B}_{\mathrm{crys}}$ and $\varphi(t^{r})=p^{r}t^{r}$, we may also assume that $r=0$. Passing to a basis of $X_{\mathrm{pket}}$ by log perfectoid affinoids, one reduces the proof to [Brinon, §6.2.2].
@@
is exact for every $n\in\mathbb{N}$ by [LLZ, Lemma 2.1.5]. ∎
@@
$\mathbf{Syn}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}(\mathcal{L},r),\quad\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c\text{-}\infty}(\mathcal{L},r).$
@@
and, using the identification $\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)\bigr)\cong\mathrm{H}^{i}\bigl(\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\bigr)$ of Proposition 6.1,
$\rho_{\mathrm{Syn},c}^{i}(r)\colon\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)\bigr)\longrightarrow\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c\text{-}\infty}(\mathcal{L},r)\bigr)\cong\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U,\mathcal{L}(r)\bigr)\bigr)[p^{-1}].$
@@ 11 Relation with Hyodo-Kato cohomology II
We say that $\mathcal{X}$ is of relative dimension $d$ when it is the finite disjoint union of formal schemes with geometrically connected special fibers all of the same dimension $d$. If there is just one component, we say that $\mathcal{X}$ is geometrically connected of relative dimension $d$.
@@
where the morphism on the bottom line is the one in Theorem 3, the most left vertical morphism is defined in Proposition 5.1 by specializing to $Z=0$, the first top morphism is the morphism $\gamma_{\rm{an}}$ given by specializing $Z\mapsto\pi$, the isomorphism
@@
is the one in [DLLZb, Thm. 3.2.7(3)] and finally we have Fontaine’s morphism
@@
See [AI, Lemma 3.25]. Using that $\frac{dX_{2}}{X_{2}}$, $\ldots$, $\frac{dX_{a}}{X_{a}}$, $\frac{dY_{1}}{Y_{1}}$, $\ldots$, $\frac{dY_{b}}{Y_{b}}$ form a basis of generators of the logarithmic differentials $\Omega^{\log}_{\widetilde{R}_{\beta}/\mathfrak{S}}$ one computes that the de Rham complex is exact except in degree 0 where the kernel of the derivative on $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}(\mathcal{U}_{\beta})$ is ${\mathcal{O}}_{\mathfrak{S}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$. ∎
@@
Our assumption that $\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)$ is a semistable representation and [Breuil, Thm. 3.3] imply that we have a Frobenius equivariant isomorphism of $B_{\rm log}$-modules :
@@
First of all note that Fontaine’s ring $B_{\rm st}$ is a subring of $B_{\rm log}$, see Remark 7.2, and the composite with the natural map $B_{\rm log}\to B_{\rm dR}$ is Fontaine’s map of rings $B_{\rm st}\to B_{\rm dR}$. Hence, the isomorphism (7) is compatible with Fontaine’s isomorphism given by replacing $B_{\rm log}$ with $B_{\rm st}$ and $B_{\rm dR}$. In fact, $D_{?}\bigl(\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\bigr)$ is defined by taking the $G_{K}$-invariants of $\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\otimes_{\mathbb{Z}_{p}}B_{?}$ for $\rm st$} or $\rm dR$}. The de Rham comparison in [DLLZb, Thm. 3.2.7(3)] follows arguing as in Proposition 11.2 and Corollary 11.3 using the period ring $\mathcal{O}\mathbb{B}_{\rm dR}$ instead of $\mathcal{O}_{\mathfrak{S}}\mathbb{B}_{\mathrm{crys}}$ and the logarithmic de Rham cohomology instead of the log-crystalline one. The conclusion of the Corollary follows as by assumption (iii) in Definition 9 the two comparisons, crystalline and de Rham, are compatible via the natural map of sheaves $\mathcal{O}_{\mathfrak{S}}\mathbb{B}_{\mathrm{crys}}\to\mathcal{O}\mathbb{B}_{\rm dR}$.
@@ 11.2 The ring $B_{\rm max}$
where $P_{\pi}(Z)$ is the minimal polynomial of $\pi$ ; see [AI, Lemma 3.23]. In particular, it contains $A_{\rm log}$ as a subring and $B_{\rm max}$ contains $B_{\rm log}$ as a subring.
@@
is flat. Notice that Frobenius on $\mathcal{O}_{K_{0}}[\![Z]\!]\otimes_{\mathcal{O}_{K_{0}}}A_{\mathrm{inf}}(\mathbb{C}_{p})$ defines a Frobenius $\varphi$ on $A_{\rm max}$ and $B_{\rm max}$, which is Frobenius on $A_{\mathrm{inf}}(\mathbb{C}_{p})$ and sends $Z\mapsto Z^{p}$ and $u\mapsto u^{p}$. Since $A_{\rm log}\cong A_{\mathrm{cris}}(\mathbb{C}_{p})\{\langle u-1\rangle\}\cong A_{\mathrm{inf}}(\mathbb{C}_{p})\{\langle P_{\pi}([\overline{\pi}]),u-1\rangle\}$ (see [AI, §2.1.1]), we have $\varphi(A_{\rm max})\subset A_{\rm log}\subset A_{\rm max}$ and $\varphi(B_{\rm max})\subset B_{\rm log}\subset B_{\rm max}$.

### v2

@@ 10 Comparison with étale cohomology
We have $\mathrm{H}^{i}\bigl(X_{\mathrm{pket}},\widehat{\mathcal{L}}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}_{\mathrm{et}}\bigl(U,\mathcal{L}(r)\bigr)$. Similarly we have $\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U,\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\mathrm{pket}},j_{!}\bigl(\widehat{\mathcal{L}}(r)\bigr)\bigr)$ where, following [LLZ23, Def. 2.3.2], we define $\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U,\mathcal{L}(r)\bigr):=\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},j_{!}\bigl(\mathcal{L}(r)\bigr)\bigr)$.
@@
We start with the first claim. First we have isomorphisms if we replace $\mathcal{L}$ with $\mathcal{L}/p^{n}\mathcal{L}$ for the isomorphisms $\mathrm{H}^{i}\bigl(X_{\mathrm{pket}},\widehat{\mathcal{L}}(r)/p^{n}\widehat{\mathcal{L}}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ by [DLL+02, Prop. 5.1.7] and $\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}_{\mathrm{et}}\bigl(U,\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ by [DLL+02, Cor. 4.6.7]. They are also finite modules ; thanks to [DLL+02, Thm. 6.2.1] the groups $\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ are finite and vanish for $i>2\mathrm{dim}X$. Using the Hochschild-Serre spectral sequence this implies that also $\mathrm{H}^{i}\bigl(X_{\mathrm{ket}},\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)$ is finite for every $i$. In particular, the inverse system for varying $n$ is Miitag-Leffler and the result follows.
For the second claim one argues similarly, using the finiteness and the vanishing for $i>2\mathrm{dim}X$ of the group $\mathrm{H}^{i}_{c}\bigl(X_{\overline{K},\mathrm{ket}},j_{!}\bigl(\mathcal{L}(r)/p^{n}\mathcal{L}(r)\bigr)\bigr)$ which is the content of [LLZ23, Thm. 2.3.5]. The isomorphism $\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\cong\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},j_{\ast}(\mathcal{L}(r))\bigr)$ and last statements for the de Rham period sheaves follow from [DLL+23, Lemma 3.6.1] using [DLL+02, Thm 6.2.1 & Cor. 6.3.4]. For the crystalline statement it suffices to show that the natural map $\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},j_{\ast}(\mathcal{L}(r))\bigr)\otimes A_{\mathrm{crys}}\to\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{A}_{\mathrm{crys}}\bigr)$ is an isomorphism. Since $p$ is a regular element in $\mathbb{A}_{\mathrm{crys}}$ by [AI12, Cor. 3.27], we are reduced to prove that the natural map $\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},j_{\ast}(\mathcal{L}(r))\bigr)\otimes A_{\mathrm{crys}}/pA_{\mathrm{crys}}\to\mathrm{H}^{i}\bigl(X_{\overline{K},\mathrm{pket}},\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{A}_{\mathrm{crys}}/p\mathbb{A}_{\mathrm{crys}}\bigr)$ is an isomorphism. As $\mathbb{A}_{\mathrm{crys}}/p\mathbb{A}_{\mathrm{crys}}$ is a free module over $\mathcal{O}_{X_{\mathrm{pket}}}^{+}/p\mathcal{O}_{X_{\mathrm{pket}}}^{+}$ and $A_{\mathrm{crys}}/pA_{\mathrm{crys}}$ is a free $\mathcal{O}_{\overline{K}}/p\mathcal{O}_{\overline{K}}$ with the same basis by [AI12, Cor. 3.24], one is reduced to the primitive comparison theorem [DLL+02, Thm 6.2.1].
@@
$0\longrightarrow j_{!}\bigl(\widehat{\mathcal{L}}(r)|_{U}\bigr)[p^{-1}]\longrightarrow\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}\stackrel{{\scriptstyle\iota\circ(1-\frac{\varphi}{p^{r}}),\iota)}}{{\longrightarrow}}\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}\oplus\widehat{\mathcal{L}}\otimes_{\widehat{\mathbb{Z}}_{p}}\mathrm{Gr}^{r}\mathbb{B}_{\mathrm{dR}}^{c-\infty}\longrightarrow 0.$
@@
The first two sequences are obtained from the analogous sequence with $\mathcal{L}=\mathbb{Z}_{p}$. Hence, we can assume that $\mathcal{L}=\mathbb{Z}_{p}$. Since $\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{crys}}=t^{r}\mathrm{Fil}^{0}\mathbb{B}_{\mathrm{crys}}$ and $\varphi(t^{r})=p^{r}t^{r}$, we may also assume that $r=0$. Passing to a basis of $X_{\mathrm{pket}}$ by log perfectoid affinoids, one reduces the proof to [BRI08, §6.2.2].
@@
is exact for every $n\in\mathbb{N}$ by [LLZ23, Lemma 2.1.5]. ∎
@@
$\mathbf{Syn}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}(\mathcal{L},r),\quad\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c-\infty}(\mathcal{L},r).$
@@
and, using the identification $\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},r)\bigr)\cong\mathrm{H}^{i}\bigl(\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\bigr)$ of Proposition 6.1,
$\rho_{\mathrm{Syn},c}^{i}(r)\colon\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},r)\bigr)\longrightarrow\mathrm{H}^{i}\bigl(\mathbf{Syn}^{c-\infty}(\mathcal{L},r)\bigr)\cong\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U,\mathcal{L}(r)\bigr)\bigr)[p^{-1}].$
@@ 11 Relation with Hyodo-Kato cohomology II
We say that $\mathcal{X}$ is of relative dimension $d$ when it is the finite disjoint union of formal schemes with geometrically connected generic and special fibers all of the same dimension $d$. If there is just one component, we say that $\mathcal{X}$ is geometrically connected of relative dimension $d$.
@@
where the morphism on the bottom line is the one in Theorem 3, the most left vertical morphism is defined in Proposition 5.1 by specializing to $Z=0$, the first top morphism is the morphism $\gamma^{\rm{an}}$ given by specializing $Z\mapsto\pi$, the isomorphism
@@
is the one in [DLL+23, Thm. 3.2.7(3)] and finally we have Fontaine’s morphism
@@
See [AI12, Lemma 3.25]. Using that $\frac{dX_{2}}{X_{2}}$, $\ldots$, $\frac{dX_{a}}{X_{a}}$, $\frac{dY_{1}}{Y_{1}}$, $\ldots$, $\frac{dY_{b}}{Y_{b}}$ form a basis of generators of the logarithmic differentials $\Omega^{\log}_{\widetilde{R}_{\beta}/\mathfrak{S}}$ one computes that the de Rham complex is exact except in degree 0 where the kernel of the derivative on $\mathcal{O}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}(\mathcal{U}_{\beta})$ is ${\mathcal{O}}_{\mathfrak{S}}\mathbb{A}_{\mathrm{crys}}\bigl(\widehat{\overline{R}}_{\beta}\bigr)$. ∎
@@
Our assumption that $\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)$ is a semistable representation and [BRE97, Thm. 3.3] imply that we have a Frobenius equivariant isomorphism of $B_{\rm log}$-modules :
@@
First of all note that Fontaine’s ring $B_{\rm st}$ is a subring of $B_{\rm log}$, see Remark 7.2, and the composite with the natural map $B_{\rm log}\to B_{\rm dR}$ is Fontaine’s map of rings $B_{\rm st}\to B_{\rm dR}$. Hence, the isomorphism (7) is compatible with Fontaine’s isomorphism given by replacing $B_{\rm log}$ with $B_{\rm st}$ and $B_{\rm dR}$. In fact, $D_{?}\bigl(\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\bigr)$ is defined by taking the $G_{K}$-invariants of $\mathrm{H}^{i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\otimes_{\mathbb{Z}_{p}}B_{?}$ for $\rm st$} or $\rm dR$}. The de Rham comparison in [DLL+23, Thm. 3.2.7(3)] follows arguing as in Proposition 11.2 and Corollary 11.3 using the period ring $\mathcal{O}\mathbb{B}_{\rm dR}$ instead of $\mathcal{O}_{\mathfrak{S}}\mathbb{B}_{\mathrm{crys}}$ and the logarithmic de Rham cohomology instead of the log-crystalline one. The conclusion of the Corollary follows as by assumption (iii) in Definition 9 the two comparisons, crystalline and de Rham, are compatible via the natural map of sheaves $\mathcal{O}_{\mathfrak{S}}\mathbb{B}_{\mathrm{crys}}\to\mathcal{O}\mathbb{B}_{\rm dR}$.
@@ 11.2 The ring $B_{\rm max}$
where $P_{\pi}(Z)$ is the minimal polynomial of $\pi$ ; see [AI12, Lemma 3.23]. In particular, it contains $A_{\rm log}$ as a subring and $B_{\rm max}$ contains $B_{\rm log}$ as a subring.
@@
is flat. Notice that Frobenius on $\mathcal{O}_{K_{0}}[\![Z]\!]\otimes_{\mathcal{O}_{K_{0}}}A_{\mathrm{inf}}(\mathbb{C}_{p})$ defines a Frobenius $\varphi$ on $A_{\rm max}$ and $B_{\rm max}$, which is Frobenius on $A_{\mathrm{inf}}(\mathbb{C}_{p})$ and sends $Z\mapsto Z^{p}$ and $u\mapsto u^{p}$. Since $A_{\rm log}\cong A_{\mathrm{cris}}(\mathbb{C}_{p})\{\langle u-1\rangle\}\cong A_{\mathrm{inf}}(\mathbb{C}_{p})\{\langle P_{\pi}([\overline{\pi}]),u-1\rangle\}$ (see [AI12, §2.1.1]), we have $\varphi(A_{\rm max})\subset A_{\rm log}\subset A_{\rm max}$ and $\varphi(B_{\rm max})\subset B_{\rm log}\subset B_{\rm max}$.